OpenStudy (anonymous):

Determine whether each series is arithmetic or geometric. Then find the sum of the given term. a) 2 + 4 + 8 + 16 + . . . ; 10th term. b) 2 + 4 + 6 + 8 + . . . ; 20th term

OpenStudy (anonymous):

For the first one I got Geometric Sequence r=2 N=10, a1=2, r=2 An=a1*rn-1 A10=2*210-1 A10=2*512 A10=1024 and the second one I got Arithmetic Sequence d=2 N=2, a1=2, d=2 An=a1+(n-1)d A20=2+(20-1)2 A20=2+(19)2 A20=2+38 A20= 40

OpenStudy (anonymous):

Can any one tell me if these are both correct and if not why?

OpenStudy (anonymous):

i ll try..

OpenStudy (anonymous):

@boocoo I can vouch for you for the second one, it is indeed an Arithmetic sequence and I got A20 as 40 too. Oddly haven't done geometric sequences, so can't vouch for you there.

OpenStudy (anonymous):

The formula 4 a geometric sequence is a1 (r ^n-1) rite? i cnt remember..

OpenStudy (anonymous):

yes that is the right formula.

OpenStudy (anonymous):

2 (2^10-1) 2(2^9) 1024 yeah ur first 1 is rite

OpenStudy (anonymous):

ok thanks :)